OpenStudy (anonymous):

Which property of equality justifies the following statement? If -8 +k=-5 then k=3

OpenStudy (rsadhvika):

-8 + k = -5 by addition law of equality, adding 8 both sides -8 + 8 +k = -5+8 k = 3